What volume (in ml) of 4.35 M HCl solution is needed to react with 18.5g of CA(OH)2
1
Expert's answer
2020-11-26T05:22:59-0500

first calculate the moles of calcium hydroxide(18.5/74.1)=0.2497moles

the equation of reaction gives a mole ratio between calcium hydroxide and hydrochloric acid as 1:2

therefor to find the moles used for HCl=(0.2497 × 2)=0.499moles

to find this vollume,(0.499 × 1000) ÷ 4.35=114.7ml